# Elementary commutator method for the Dirac equation with long-range perturbations

### v1

@@ 1.1.1 Introduction
In this paper we investigate the Dirac equation in the $d$-dimensional configuration space $\mathbb{R}^{d}$ with $d\in\mathbb{N}=\{1,2,\ldots\}$. We write it in the form of evolution equation as
@@ 1.2.1 Rellich’s theorem
    The function space $\mathcal{B}^{*}_{0}$ is optimal in the sense that there actually exists a non-trivial solution to (1.10) in $\mathcal{B}^{*}$, see Remark 1.14 for the construction. The largest function space so far is $L^{2}_{-1/2}(\mathbb{R}^{d})$ for $d=3$ due to Vogelsang [31], as far as the authors are aware of.
@@ 1.2.3 Radiation condition bounds and applications
Remark 1.14.
Using Corollary 1.13, we can construct a non-trivial solution to (1.10). In fact, for any $\psi\in L^{2}_{1/2+\kappa}$ with $\kappa\in(0,\rho/2)$ we set
$\phi=R_{+}(\lambda)\psi-R_{-}(\lambda)\psi.$
Then it satisfies (1.10) due to 1 of Corollary 1.13, and is non-trivial due to 2 of Corollary 1.13 and Theorem 1.3.
@@ 2 Proof of Rellich’s theorem
2.1 Reduction to three propositions
In this section we prove Theorem 1.3, following the scheme of Ito–Skibsted [16]. We split the proof into three steps. Obviously, Theorem 1.3 is a consequence of the following propositions. Throughout the section we assume Assumption 1.1.
@@ 2.1 Reduction to three propositions
Then there exists $\mu>0$ such that $\mathrm{e}^{\mu f}\phi\in\mathcal{B}_{0}^{*}$.
Proposition 2.2.
  1. 1.
    $(H-\lambda)\phi=0$ in the distributional sense,
  2. 2.
    $\mathrm{e}^{\kappa f}\phi\in\mathcal{B}_{0}^{*}$ for some $\kappa>0$.
Then for any $\mu>0$ one has $\mathrm{e}^{(\kappa+\mu)f}\phi\in\mathcal{B}^{*}_{0}$.
@@
We will prove these propositions in Sections 2.3, 2.4 and 2.5, respectively, after short preliminaries in Section $2.$
@@ 2.2.1 Conjugate operator
Here we present formulas, to be repeatedly referred to in this and later sections, involving a general weight function. The proofs of Propositions 2.1, 2.2 and 2.3 depend on a commutator method sharing conjugate operators of the same form, however, with different weight functions.
@@ 2.2.2 Weight functions for Rellich’s theorem
$0<\int_{0}^{f}(1+\tau/2^{\nu})^{-1-\delta}\,\mathrm{d}\tau\leq\min\{f,\delta 2^{\nu}\}\ \ \text{for any }\nu\in\mathbb{N}_{0},$
@@
In particular, noting $2^{-\nu}(1+f/2^{\nu})^{-1}\leq f^{-1}$, we can bound for any $k\in\mathbb{N}$
$|(\theta-2\kappa f)^{(k)}|\leq C_{\delta,k}f^{-k}(\theta^{\prime}-2\kappa).$
@@ 2.3 A priori exponential decay estimate
Let $\lambda\gtrless m_{\pm}$, respectively, and let $\kappa=0$ in the definition (2.6) of $\Theta$. Then there exist $\mu,c,C>0$ and $a_{0}\in\mathbb{N}_{0}$ such that uniformly in $b>a\geq a_{0}$ and $\nu\in\mathbb{N}_{0}$
@@
respectively, where $\gamma=\gamma_{a,b}$ is a certain function satisfying $\mathop{\mathrm{supp}}\gamma\subset\mathop{\mathrm{supp}}\chi_{a,b}$ and $|\gamma|\leq C_{a,b}$.
@@
    We can see that the first term on the right-hand side is exactly from the first and second term of Lemma 2.4 with $\Theta$ of the form (2.6).
@@
    Let us outline the deduction of Proposition 2.1 from Lemma 2.8: Take the expectation of the above inequality on the eigenstate $\phi$, and the left-hand side and the last term on the right-hand side vanish; Let $b\to\infty$, and a contribution from $\chi_{b-1,b+1}^{2}f^{-1}\mathrm{e}^{\theta}$ vanishes by definition of $\mathcal{B}^{*}_{0}$; Then let $\nu\to\infty$, which replaces the Yosida approximation $\theta$ by $f$. Proposition 2.1 thus follows.
@@
Let us only discuss the upper sign since the lower one can be done by the same manner. Fix $\lambda>m_{\pm}$ and $\kappa=0$ as in the assertion. For the moment we merely choose $a_{0}\in\mathbb{N}_{0}$ such that for some $c_{1}>0$
$\lambda-q_{0}\geq c_{1}\ \ \text{uniformly in }\{x\in\mathbb{R}^{d};\ |x|\geq 2^{a_{0}}\}.$ (2.9)
Then, except for the last part of the proof, the below estimates are all uniform in $\mu\in(0,1]$, $b>a\geq a_{0}$ and $\nu\in\mathbb{N}_{0}$, and thus until then $c_{*},C_{*}>0$ are independent of them. Only in the very last step we shall restrict ranges of all these parameters to verify the assertion, where we may also retake $a_{0}\in\mathbb{N}_{0}$ larger if necessary.
@@
$\begin{split}Q&=\bigl(\chi_{a,b}f^{-1}\theta^{\prime}+\chi_{a,b}f^{-1-\rho}+|\chi_{a,b}^{\prime}|\bigr)\mathrm{e}^{\theta}\\ &\phantom{{}={}}{}+p_{j}\bigl(\chi_{a,b}f^{-1}\theta^{\prime}+\chi_{a,b}f^{-1-\rho}+|\chi_{a,b}^{\prime}|\bigr)\mathrm{e}^{\theta}p_{j}+(H-\lambda)f^{1+\rho}\Theta(H-\lambda),\end{split}$ (2.10)
which we prove at last is negligible.
@@
$\begin{split}&\mathop{\mathrm{Re}}\bigl(\alpha_{f}\theta^{\prime}\Theta p_{f}\bigr)+\mathop{\mathrm{Re}}\bigl(\alpha_{j}f^{-1}\Theta\ell_{jk}p_{k}\bigr)\\ &=\tfrac{1}{2}\mathop{\mathrm{Re}}\bigl(\alpha_{f}(\theta^{\prime}+f^{-1})\Theta p_{f}\bigr)-\tfrac{1}{2}\mathop{\mathrm{Re}}\bigl(\alpha_{f}(f^{-1}-\theta^{\prime})\Theta p_{f}\bigr)\\ &\phantom{{}={}}+\tfrac{1}{2}\mathop{\mathrm{Re}}\bigl(\alpha_{j}(\theta^{\prime}+f^{-1})\Theta\ell_{jk}p_{k}\bigr)+\tfrac{1}{2}\mathop{\mathrm{Re}}\bigl(\alpha_{j}(f^{-1}-\theta^{\prime})\Theta\ell_{jk}p_{k}\bigr)\\ &=\tfrac{1}{2}\mathop{\mathrm{Re}}\bigl(\alpha_{j}(\theta^{\prime}+f^{-1})\Theta p_{j}\bigr)-\tfrac{1}{2}\mathop{\mathrm{Re}}\bigl(\alpha_{j}(f^{-1}-\theta^{\prime})\Theta h_{jk}p_{k}\bigr)\\ &=\tfrac{1}{2}(\lambda-q)(\theta^{\prime}+f^{-1})\Theta+\tfrac{1}{2}\mathop{\mathrm{Re}}\bigl((\theta^{\prime}+f^{-1})\Theta(H-\lambda)\bigr)\\ &\phantom{{}={}}-\tfrac{1}{2}\mathop{\mathrm{Re}}\bigl(\mathrm{e}^{\theta/2}\chi_{a,b}(f^{-1}-\theta^{\prime})\alpha_{j}h_{jk}p_{k}\mathrm{e}^{\theta/2}\bigr)\\ &\geq\tfrac{1}{2}(\lambda-q_{0}-q_{1})(\theta^{\prime}+f^{-1})\Theta-\tfrac{1}{4}(\lambda-q_{0})|f^{-1}-\theta^{\prime}|\Theta\\ &\phantom{{}={}}-\tfrac{1}{4}\mathrm{e}^{\theta/2}p_{i}h_{il}\alpha_{l}\chi_{a,b}(\lambda-q_{0})^{-1}|f^{-1}-\theta^{\prime}|\alpha_{j}h_{jk}p_{k}\mathrm{e}^{\theta/2}-C_{2}Q,\end{split}$ (2.12)
@@
cf. (1.6), and we proceed with the third term of (2.12) by using (1.1), (2.3) and the Cauchy–Schwarz inequality as
$\begin{split}&-\tfrac{1}{4}\mathrm{e}^{\theta/2}p_{i}h_{il}\alpha_{l}\chi_{a,b}(\lambda-q_{0})^{-1}|f^{-1}-\theta^{\prime}|\alpha_{j}h_{jk}p_{k}\mathrm{e}^{\theta/2}\\ &\geq-\tfrac{1}{4}\mathrm{e}^{\theta/2}p_{i}h_{il}\alpha_{l}\Xi\alpha_{j}h_{jk}p_{k}\mathrm{e}^{\theta/2}\\ &=-\tfrac{1}{4}\mathop{\mathrm{Re}}\bigl(\alpha_{l}\alpha_{j}\widetilde{\Xi}\mathrm{e}^{\theta/2}h_{il}h_{jk}p_{i}p_{k}\mathrm{e}^{\theta/2}\bigr)-\tfrac{1}{4}\mathop{\mathrm{Im}}\bigl(\alpha_{l}\mathrm{e}^{\theta/2}(\partial_{i}h_{il}\Xi h_{jk})\alpha_{j}p_{k}\mathrm{e}^{\theta/2}\bigr)\\ &=-\tfrac{1}{4}\mathop{\mathrm{Re}}\bigl(\widetilde{\Xi}\mathrm{e}^{\theta/2}h_{ij}h_{jk}p_{i}p_{k}\mathrm{e}^{\theta/2}\bigr)-\tfrac{1}{4}\mathop{\mathrm{Im}}\bigl(\alpha_{l}(\partial_{i}h_{il}\Xi h_{jk})\alpha_{j}\mathrm{e}^{\theta}p_{k}\bigr)\\ &\phantom{{}={}}+\tfrac{1}{8}\mathop{\mathrm{Re}}\bigl(\alpha_{l}(\partial_{i}h_{il}\Xi h_{jk})\alpha_{j}(\partial_{k}f)\theta^{\prime}\mathrm{e}^{\theta}\bigr)\\ &\geq-\tfrac{1}{4}\mathop{\mathrm{Re}}\bigl(\widetilde{\Xi}\mathrm{e}^{\theta/2}\delta_{ik}p_{i}p_{k}\mathrm{e}^{\theta/2}\bigr)-C_{3}Q.\end{split}$ (2.14)
Using (1.1) and the Cauchy–Schwarz inequality again, and noting $\alpha_{f}^{2}=1$, we continue (2.14) as
@@
$\begin{split}&-\tfrac{1}{4}\mathrm{e}^{\theta/2}p_{i}h_{il}\alpha_{l}\chi_{a,b}(\lambda-q_{0})^{-1}|f^{-1}-\theta^{\prime}|\alpha_{j}h_{jk}p_{k}\mathrm{e}^{\theta/2}\\ &\geq-\tfrac{1}{4}\mathop{\mathrm{Re}}\bigl(\widetilde{\Xi}\mathrm{e}^{\theta/2}\alpha_{i}\alpha_{k}p_{i}p_{k}\mathrm{e}^{\theta/2}\bigr)-C_{3}Q\\ &=-\tfrac{1}{4}\mathrm{e}^{\theta/2}p_{i}\alpha_{i}\Xi\alpha_{k}p_{k}\mathrm{e}^{\theta/2}+\tfrac{1}{4}\mathop{\mathrm{Im}}\bigl(\mathrm{e}^{\theta/2}\alpha_{i}(\partial_{i}\Xi)\alpha_{k}p_{k}\mathrm{e}^{\theta/2}\bigr)-C_{3}Q\\ &=-\tfrac{1}{4}\bigl(p_{i}\alpha_{i}+\tfrac{\mathrm{i}}{2}\alpha_{f}\theta^{\prime}\bigr)\Xi\mathrm{e}^{\theta}\bigl(\alpha_{k}p_{k}-\tfrac{\mathrm{i}}{2}\alpha_{f}\theta^{\prime}\bigr)+\tfrac{1}{4}\mathop{\mathrm{Im}}\bigl(\alpha_{i}(\partial_{i}\Xi)\mathrm{e}^{\theta}\alpha_{k}p_{k}\bigr)\\ &\phantom{{}={}}-\tfrac{1}{8}\mathop{\mathrm{Re}}\bigl(\alpha_{i}\alpha_{f}(\partial_{i}\Xi)\theta^{\prime}\mathrm{e}^{\theta}\bigr)-C_{3}Q\\ &\geq-\tfrac{1}{4}p_{i}\alpha_{i}\Xi\mathrm{e}^{\theta}\alpha_{k}p_{k}+\tfrac{1}{8}\alpha_{i}\alpha_{k}\bigl(\partial_{i}\Xi\theta^{\prime}(\partial_{k}f)\mathrm{e}^{\theta}\bigr)-\tfrac{1}{16}\Xi\theta^{\prime 2}\mathrm{e}^{\theta}-C_{4}Q\\ &\geq-\tfrac{1}{4}(\lambda-q+H-\lambda)\Xi\mathrm{e}^{\theta}(\lambda-q+H-\lambda)+\tfrac{1}{16}\Xi\theta^{\prime 2}\mathrm{e}^{\theta}-C_{5}Q\\ &\geq-\tfrac{1}{4}(\lambda-q)\Xi\mathrm{e}^{\theta}(\lambda-q)-C_{6}Q\\ &\geq-\tfrac{1}{4}(\lambda-q_{0}-2q_{1})|f^{-1}-\theta^{\prime}|\Theta-C_{7}Q.\end{split}$ (2.15)
Hence by (2.11), (2.12) and (2.15) we obtain
@@
$\begin{split}\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-\lambda)\bigr)&\geq(c_{1}-q_{1})\min\{f^{-1},\theta^{\prime}\}\Theta-C_{8}Q.\end{split}$ (2.16)
Next we discuss $Q$. Here we claim

### v2

@@ Abstract
Keywords: Dirac operator, spectral theory, eigenfunctions, resolvent estimates
Mathematics Subject Classification 2020: 35P10, 35Q41, 81Q15
@@ 1.1.1 Introduction
In this paper we investigate the Dirac equation on the Euclidean space $\mathbb{R}^{d}$ of dimension $d\in\mathbb{N}=\{1,2,\ldots\}$. We write it in the form of evolution equation as
@@ 1.2.1 Rellich’s theorem
    The function space $\mathcal{B}^{*}_{0}$ is optimal since we can in fact construct a non-trivial solution to (1.10) in $\mathcal{B}^{*}$ by using a WKB-type approximation and Corollary 1.13 below, see, e.g., [18, Proposition 4.18]. So far the largest space where the generalized eigenfunctions are absent is $L^{2}_{-1/2}(\mathbb{R}^{d})$ for $d=3$ due to Vogelsang [31], as far as the authors are aware of.
@@ 2 Proof of Rellich’s theorem
2.1 Reduction to two propositions
In this section we prove Theorem 1.3, following the scheme of Ito–Skibsted [16]. We split the proof into two steps. Obviously, Theorem 1.3 is a consequence of the following propositions. Throughout the section we assume Assumption 1.1.
@@ 2.1 Reduction to two propositions
Then for any $\kappa>0$ one has $\mathrm{e}^{\kappa f}\phi\in\mathcal{B}_{0}^{*}$.
@@
We will prove these propositions in Sections 2.3 and 2.4, respectively, after short preliminaries in Section $2.$
@@ 2.2.1 Conjugate operator
Here we present formulas, to be repeatedly referred to in this and later sections, involving a general weight function. The proofs of Propositions 2.1 and 2.2 depend on a commutator method sharing conjugate operators of the same form, however, with different weight functions.
@@ 2.2.2 Weight functions for Rellich’s theorem
$0<\int_{0}^{f}(1+\tau/2^{\nu})^{-1-\delta}\,\mathrm{d}\tau\leq\min\{f,\delta^{-1}2^{\nu}\}\ \ \text{for any }\nu\in\mathbb{N}_{0},$
@@
In particular, noting $2^{-\nu}(1+f/2^{\nu})^{-1}\leq f^{-1}$, we can bound for any $k=2,3,\ldots$
$|(\theta-2\kappa f)^{\prime}|\leq C_{\delta,1}(\theta^{\prime}-2\kappa),\qquad|(\theta-2\kappa f)^{(k)}|\leq C_{\delta,k}f^{2-k}|\theta^{\prime\prime}|.$
@@ 2.3 A priori exponential decay estimate
Let $\lambda\gtrless m_{\pm}$, respectively, and fix any $\kappa_{0}\geq 0$. Then there exist $\mu,c,C>0$ and $a_{0}\in\mathbb{N}_{0}$ such that uniformly in $\kappa\in[0,\kappa_{0}]$, $b>a\geq a_{0}$ and $\nu\in\mathbb{N}_{0}$
@@
respectively, where $\gamma=\gamma_{a,b}$ is a certain function, independent of $\kappa$ and $\nu$, satisfying $\mathop{\mathrm{supp}}\gamma\subset\mathop{\mathrm{supp}}\chi_{a,b}$ and $|\gamma|\leq C_{a,b}$.
@@
    The first term on the right-hand side is from the first and second term of Lemma 2.3 with $\Theta$ of the form (2.6). It is exactly the minimum of the coefficients
@@
    Let us outline the deduction of Proposition 2.1 from Lemma 2.7: Assume
    $\kappa_{0}=\sup\{\kappa\geq 0;\ \mathrm{e}^{\kappa f}\phi\in\mathcal{B}^{*}_{0}\}<\infty,$ (2.9)
    let $\mu>0$ be as above, and let $\kappa=0$ if $\kappa_{0}=0$, and $\kappa\in[0,\kappa_{0})$ with $\kappa+\mu>\kappa_{0}$ otherwise; Take the expectation of the above inequality on the eigenstate $\phi$, and the left-hand side and the last term on the right-hand side vanish; Let $b\to\infty$, and a contribution from $\chi_{b-1,b+1}^{2}f^{-1}\mathrm{e}^{\theta}$ vanishes by definition of $\mathcal{B}^{*}_{0}$; Let $\nu\to\infty$, which replaces the Yosida approximation $\theta$ by $f$; Then we obtain $\mathrm{e}^{(\kappa+\mu)f}\phi\in\mathcal{B}^{*}_{0}$, which contradicts the assumption. Proposition 2.1 thus follows. The second term on the right-hand side of Lemma 2.3 may be seen as a boundary contribution, as in Gauss’s divergence theorem.
@@
Let us only discuss the upper sign since the lower one can be done by the same manner. Fix $\lambda>m_{+}$ and $\kappa_{0}\geq 0$ as in the assertion. For the moment we merely choose $a_{0}\in\mathbb{N}_{0}$ such that for some $c_{1}>0$
$\min\{\lambda-q_{0},\lambda-q_{0}-q_{1}\}\geq c_{1},\ \ \text{uniformly in }\{x\in\mathbb{R}^{d};\ |x|\geq 2^{a_{0}}\}.$ (2.10)
Then, except for the last part of the proof, the below estimates are all uniform in $\kappa\in[0,\kappa_{0}]$, $\mu\in(0,1]$, $b>a\geq a_{0}$ and $\nu\in\mathbb{N}_{0}$, and thus until then $c_{*},C_{*}>0$ are independent of them. Only in the very last step we shall restrict ranges of all these parameters to verify the assertion, where we may also retake $a_{0}\in\mathbb{N}_{0}$ larger if necessary.
@@
$\begin{split}Q&=\bigl(\chi_{a,b}|\theta^{\prime\prime}|+\chi_{a,b}f^{-1-\rho}+|\chi_{a,b}^{\prime}|\bigr)\mathrm{e}^{\theta}\\ &\phantom{{}={}}{}+p_{j}\bigl(\chi_{a,b}|\theta^{\prime\prime}|+\chi_{a,b}f^{-1-\rho}+|\chi_{a,b}^{\prime}|\bigr)\mathrm{e}^{\theta}p_{j}+(H-\lambda)f^{1+\rho}\Theta(H-\lambda).\end{split}$ (2.11)
We will later prove that it is in fact negligible.
@@
$\begin{split}&\mathop{\mathrm{Re}}\bigl(\alpha_{f}\theta^{\prime}\Theta p_{f}\bigr)+\mathop{\mathrm{Re}}\bigl(\alpha_{j}f^{-1}\Theta\ell_{jk}p_{k}\bigr)\\ &=\tfrac{1}{2}\mathop{\mathrm{Re}}\bigl(\alpha_{f}(\theta^{\prime}+f^{-1})\Theta p_{f}\bigr)-\tfrac{1}{2}\mathop{\mathrm{Re}}\bigl(\alpha_{f}(f^{-1}-\theta^{\prime})\Theta p_{f}\bigr)\\ &\phantom{{}={}}+\tfrac{1}{2}\mathop{\mathrm{Re}}\bigl(\alpha_{j}(\theta^{\prime}+f^{-1})\Theta\ell_{jk}p_{k}\bigr)+\tfrac{1}{2}\mathop{\mathrm{Re}}\bigl(\alpha_{j}(f^{-1}-\theta^{\prime})\Theta\ell_{jk}p_{k}\bigr)\\ &=\tfrac{1}{2}\mathop{\mathrm{Re}}\bigl(\alpha_{j}(\theta^{\prime}+f^{-1})\Theta p_{j}\bigr)-\tfrac{1}{2}\mathop{\mathrm{Re}}\bigl(\alpha_{j}(f^{-1}-\theta^{\prime})\Theta h_{jk}p_{k}\bigr)\\ &=\tfrac{1}{2}(\lambda-q)(\theta^{\prime}+f^{-1})\Theta+\tfrac{1}{2}\mathop{\mathrm{Re}}\bigl((\theta^{\prime}+f^{-1})\Theta(H-\lambda)\bigr)\\ &\phantom{{}={}}-\tfrac{1}{2}\mathop{\mathrm{Re}}\bigl(\chi_{a,b}(f^{-1}-\theta^{\prime})f^{2-d}\alpha_{j}h_{jk}p_{k}f^{d-2}\mathrm{e}^{\theta}\bigr)\\ &\geq\tfrac{1}{2}(\lambda-q_{0}-q_{1})(\theta^{\prime}+f^{-1})\Theta-\tfrac{1}{4}(\lambda-q_{0})|f^{-1}-\theta^{\prime}|\Theta\\ &\phantom{{}={}}-\tfrac{1}{4}\mathrm{e}^{\theta}f^{d-2}p_{i}h_{il}\alpha_{l}\chi_{a,b}(\lambda-q_{0})^{-1}|f^{-1}-\theta^{\prime}|f^{4-2d}\mathrm{e}^{-\theta}\alpha_{j}h_{jk}p_{k}f^{d-2}\mathrm{e}^{\theta}\\ &\phantom{{}={}}-C_{2}Q,\end{split}$ (2.13)
@@
cf. (1.6), and we proceed with the third term of (2.13) by using (2.14), (2.3) and $\alpha_{f}^{2}=1$ as
@@
$\begin{split}&-\tfrac{1}{4}\mathrm{e}^{\theta}f^{d-2}p_{i}h_{il}\alpha_{l}\chi_{a,b}(\lambda-q_{0})^{-1}|f^{-1}-\theta^{\prime}|f^{4-2d}\mathrm{e}^{-\theta}\alpha_{j}h_{jk}p_{k}f^{d-2}\mathrm{e}^{\theta}\\ &\geq-\tfrac{1}{4}\mathrm{e}^{\theta}f^{d-2}p_{i}h_{il}\alpha_{l}\Xi f^{4-2d}\mathrm{e}^{-\theta}\alpha_{j}h_{jk}p_{k}f^{d-2}\mathrm{e}^{\theta}\\ &=-\tfrac{1}{4}\mathrm{e}^{\theta/2}\bigl(p_{i}+\tfrac{\mathrm{i}}{2}(\partial_{i}f)\theta^{\prime}+\mathrm{i}(d-2)(\partial_{i}f)f^{-1}\bigr)h_{il}\alpha_{l}\Xi\alpha_{j}h_{jk}\\ &\phantom{{}={}}{}\cdot\bigl(p_{k}-\tfrac{\mathrm{i}}{2}(\partial_{k}f)\theta^{\prime}-\mathrm{i}(d-2)(\partial_{k}f)f^{-1}\bigr)\mathrm{e}^{\theta/2}\\ &=-\tfrac{1}{4}\mathrm{e}^{\theta/2}p_{i}h_{il}\alpha_{l}\Xi\alpha_{j}h_{jk}p_{k}\mathrm{e}^{\theta/2}\\ &\phantom{{}={}}{}+\tfrac{1}{4}\mathop{\mathrm{Im}}\bigl(\alpha_{f}\Xi(\theta^{\prime}+2(d-2)f^{-1})\mathrm{e}^{\theta/2}\alpha_{j}h_{jk}p_{k}\mathrm{e}^{\theta/2}\bigr)\\ &\phantom{{}={}}{}-\tfrac{1}{16}\widetilde{\Xi}(\theta^{\prime}+2(d-2)f^{-1})^{2}\mathrm{e}^{\theta}\\ &\geq-\tfrac{1}{4}\mathop{\mathrm{Re}}\bigl(\alpha_{l}\alpha_{j}\mathrm{e}^{\theta/2}\widetilde{\Xi}h_{il}h_{jk}p_{i}p_{k}\mathrm{e}^{\theta/2}\bigr)-\tfrac{1}{4}\mathop{\mathrm{Im}}\bigl(\alpha_{l}\alpha_{j}\mathrm{e}^{\theta/2}(\partial_{i}\widetilde{\Xi}h_{il}h_{jk})p_{k}\mathrm{e}^{\theta/2}\bigr)\\ &\phantom{{}={}}{}+\tfrac{1}{4}\mathop{\mathrm{Im}}\bigl(\widetilde{\Xi}(\theta^{\prime}+2(d-2)f^{-1})\mathrm{e}^{\theta/2}p_{f}\mathrm{e}^{\theta/2}\bigr)\\ &\phantom{{}={}}{}-\tfrac{1}{4}\mathop{\mathrm{Im}}\bigl(\alpha_{f}\Xi(\theta^{\prime}+2(d-2)f^{-1})\mathrm{e}^{\theta/2}\alpha_{j}\ell_{jk}p_{k}\mathrm{e}^{\theta/2}\bigr)\\ &\phantom{{}={}}{}-\tfrac{1}{16}\widetilde{\Xi}\theta^{\prime 2}\mathrm{e}^{\theta}-\tfrac{d-2}{4}\widetilde{\Xi}f^{-1}\theta^{\prime}\mathrm{e}^{\theta}-C_{3}Q.\end{split}$ (2.15)
We continue to compute (2.15). Note that by (2.14), (2.3) and (2.4)
$\partial_{i}h_{jk}=2f^{-1}\ell_{ij}(\partial_{k}f)+2f^{-1}(\partial_{j}f)\ell_{ik},\quad(\Delta f)=(d-1)f^{-1}.$ (2.16)
Then, after some computations employing (1.1), Assumption 1.1, the Cauchy–Schwarz inequality, (2.16) and $\ell_{ij}\alpha_{i}\alpha_{j}=(d-1)$ along with (2.3), (2.4), (2.14) and $\alpha_{f}^{2}=1$ as above, we obtain
@@
$\begin{split}&-\tfrac{1}{4}\mathrm{e}^{\theta}f^{d-2}p_{i}h_{il}\alpha_{l}\chi_{a,b}(\lambda-q_{0})^{-1}|f^{-1}-\theta^{\prime}|f^{4-2d}\mathrm{e}^{-\theta}\alpha_{j}h_{jk}p_{k}f^{d-2}\mathrm{e}^{\theta}\\ &\geq-\tfrac{1}{4}\mathop{\mathrm{Re}}\bigl(\mathrm{e}^{\theta/2}\widetilde{\Xi}h_{ij}h_{jk}p_{i}p_{k}\mathrm{e}^{\theta/2}\bigr)-\tfrac{d-1}{2}\mathop{\mathrm{Im}}\bigl(\alpha_{f}\alpha_{j}\widetilde{\Xi}f^{-1}\mathrm{e}^{\theta/2}h_{jk}p_{k}\mathrm{e}^{\theta/2}\bigr)\\ &\phantom{{}={}}{}-\tfrac{1}{2}\mathop{\mathrm{Im}}\bigl(\alpha_{l}\alpha_{j}\widetilde{\Xi}f^{-1}\mathrm{e}^{\theta/2}h_{il}\ell_{ij}p_{f}\mathrm{e}^{\theta/2}\bigr)-\tfrac{1}{2}\mathop{\mathrm{Im}}\bigl(\alpha_{l}\alpha_{f}\widetilde{\Xi}f^{-1}\mathrm{e}^{\theta/2}h_{il}\ell_{ik}p_{k}\mathrm{e}^{\theta/2}\bigr)\\ &\phantom{{}={}}{}+\tfrac{d-1}{8}\widetilde{\Xi}f^{-1}\theta^{\prime}\mathrm{e}^{\theta}-\tfrac{1}{4}\mathop{\mathrm{Im}}\bigl(\alpha_{f}\Xi(\theta^{\prime}+2(d-2)f^{-1})\mathrm{e}^{\theta}\alpha_{j}\ell_{jk}p_{k}\bigr)\\ &\phantom{{}={}}{}-\tfrac{1}{16}\widetilde{\Xi}\theta^{\prime 2}\mathrm{e}^{\theta}-\tfrac{d-2}{4}\widetilde{\Xi}f^{-1}\theta^{\prime}\mathrm{e}^{\theta}-C_{4}Q\\ &\geq-\tfrac{1}{4}\mathop{\mathrm{Re}}\bigl(\mathrm{e}^{\theta/2}\widetilde{\Xi}\delta_{ik}p_{i}p_{k}\mathrm{e}^{\theta/2}\bigr)-\tfrac{1}{4}\mathop{\mathrm{Im}}\bigl(\alpha_{f}\Xi\theta^{\prime}\mathrm{e}^{\theta}\alpha_{j}\ell_{jk}p_{k}\bigr)\\ &\phantom{{}={}}{}-\tfrac{1}{16}\widetilde{\Xi}\theta^{\prime 2}\mathrm{e}^{\theta}-\tfrac{d-3}{8}\widetilde{\Xi}f^{-1}\theta^{\prime}\mathrm{e}^{\theta}-C_{5}Q.\end{split}$ (2.17)
Let us compute the first term on the right-hand side of (2.17) as follows: